find all solutions of the equation algebraically.

1/t^2 + 8/t + 15 = 0

Answer by stanbon(75887)

find all solutions of the equation algebraically.
1/t^2 + 8/t + 15 = 0
Multiply thru by t^2 to get:
1 + 8t +15t^2 = 0
Rearrange:
15t^2 + 8t + 1 = 0
(3t+1)(5t+1) = 0
t = -1/3 or t = -1/5
-------------------------
Cheers,
Stan H.